# Lindbladian dynamics with loss of quantum jumps

### v2

@@ Abstract
Introduction.— The Lindblad master equation (LME) is one of the most successful dynamics equations in the quantum open system Lindblad1976 ; Gorini1976 ; Gardiner1985 ; Daniel2020 . Its success is not only attributed to the effective description of the quantum system weakly coupling to the environment but also stems from its concise mathematical structure with good properties such as trace-preserving, trace-preserving, and complete positivity. From the perspective of quantum trajectory theory Dalibard1992 ; Dum1992 ; Gardiner1992 ; Klaus1993 ; Plenio1998 ; Daley2014 , the LME can be understood as an average of many quantum trajectories consisting of the evolution of non-Hermitian Hamiltonian (ENHH) and random quantum jumps. Due to the constraint by LME, ENHH and the quantum jumps are required to reach a balance in the average sense of all trajectories to keep some good properties like trace-preserving.
@@
Nonlinear Lindblad master equation.— Considering a system with Hamiltonian $H$ coupling to the environment by $M$ jump operators $L_{\mu}$ with strength $\gamma_{\mu}$, its density matrix $\rho$ is governed by LME:
@@
where $\mathcal{L}$ is the Liouvillian superoperator and Planck constant $\hbar$ is set to unity throughout this paper. We monitor every jump by a $\eta_{\mu}$-efficiency detector, $\eta_{\mu}\in[0,1]$, as shown in Fig. 1(a). The core concern of the present work is how $\rho$ evolves after discarding the data along with detectors clicking.
@@
we get the NLME (see Supplementary S1 SuMa for details):
@@
The simplest model for Eq. (4) is a two-level atom with monitoring its spontaneous emission, as shown in Fig. 1(b), where $J$ is the coupling strength and $\gamma$ is the spontaneous emission rate of the excited state $|e\rangle$. The detector of emission photons has a finite efficiency $\eta$. In the experiment, the observed data of atomic population is discarded when emission photons are detected simultaneously. With the remaining data, the evolution of atomic system is rebuilt by
@@
A vectorization treatment of Eq. (9) is given in Supplementary S2 SuMa . We further compare the results from this method with results from two kinds of quantum trajectory methods in Supplementary S4 SuMa to confirm the validity of NLME.
Properties and classification.— The NLME given by Eq. (4) provides a convenient method to study the postselection dynamics. Before going to the study of concrete systems, we discuss some specific features of the NLME. Firstly, we show some properties of NLME.
@@
An example in the trivial class.— Considering a homogeneous fermion chain in periodic boundary condition under $\eta$-efficiency continuous measurements of particle number in every site. The postselection dynamics is described by
@@
Postselected skin effect.— Now we study a model in the nontrivial class, which exhibits skin effect induced by postselection. Considering a $L$-site free fermion chain under open boundary condition monitored by $L+1$ detectors with unified efficiency $\eta$ as shown in Fig. 2(a), the corresponding NLME is given by
@@
where $H=J\sum_{l=1}^{L-1}a_{l}^{\dagger}a_{l+1}+h.c.$, $\gamma_{l}=\gamma$, $L_{l}=d_{l}^{\dagger}d_{l}$ with $d_{l}=(a_{l}+ia_{l+1})/\sqrt{2}$ Wang2022 for $l=1\sim L-1$, $\gamma_{0}=\gamma_{L}=\gamma/2$, $L_{0}=\hat{n}_{1}$ and $L_{L}=\hat{n}_{L}$. For simplicity, we set $\gamma\leq J=1$.
Figure 2: Postselected skin effect. (a) Illustration of the model with postselected skin effect. (b) Steady-state distributions of the half-filled chain with 50 (red star), 100 (blue circle), and 150 (black triangle) sites. The green line is the fitting function Eq. (22) with $\beta=0.04$. (c) Evolution of half-chain entanglement entropy defined by Eq. (25) in the half-filled chain with 50 (red line), 100 (blue line), and 150 (black line) sites. In (b) and (c), $\gamma=0.4$, $\eta=0.6$ and the initial state is $|1010\cdots\rangle$. The results are an average of $60$ quantum trajectories by the Gaussian state method and the Monte Carlo wave-function method SuMa with time step $\delta t=0.005$.
@@
Zeno effect versus skin effect.— The NLME of postselected skin effect formally is same as the Hatano-Nelson model in open system. The effective Hamiltonian $H_{eff}(\eta\gamma)$ in Eq. (23a) leads the non-Hermitian skin effect Yao2018 ; Kunst2018 ; CHLee2019 , while the environmental effect $\mathcal{L}_{M}$ in Eq. (23b) can be understood as continuous measurements, which induces the quantum Zeno effect Misra1997 . Considering that both of Zeno effect and skin effect can influence the entangled entropy growth, the TAEE $\overline{S}_{L}(x_{C},x_{C}+\Delta)$ will be a good indicator to reflect the two effects’ competition and cooperation. The TAEE is defined as
@@
Moreover, the distribution and scaling behavior of TAEE are significantly influenced by the postselected skin effect, both exhibiting a two-segment structure—a nearly algebraic growth followed by saturation at the end, as discussed in Supplementary S6 SuMa .
Summary and discussion.— In summary, we proposed a NLME to describe the normalized postselection dynamics of Lindblad master equations with loss of quantum jumps. It continuously interpolates between LME and ENHH by reducing quantum jumps. With the advanced analytical ability of NLME, we answer the question of how it influences the dynamics of LME by losing quantum jumps. We classify the influences into two classes. In trivial class, discarding jumps is completely equivalent to weakening the coupling strength of the environment. While in the nontrivial class, it makes a difference by breaking the balance of LME between the effective non-Hermitain Hamiltonian and quantum jumps. We show a new phenomenon called the postselected skin effect, where the steady state exhibits a special biased particle profile by discarding a few quantum jumps. Moreover, we show the coexistence and competition between Zeno effect and non-Hermitian skin effect in the postselected skin effect by capturing the characteristics of the TAEE.
@@
Acknowledgements.
Acknowledgments.— We thank Y.-P. Wang for helpful discussions on the Gaussian-state method. This work is supported by National Key Research and Development Program of China (Grant No. 2023YFA1406704 and 2021YFA1402104), the NSFC under Grants No. 12174436 and No. T2121001, and the Strategic Priority Research Program of Chinese Academy of Sciences under Grant No. XDB33000000.
SUPPLEMENTAL MATERIAL: The dynamics of Lindblad master equations with loss of quantum jumps
S1. The derivation of the nonlinear Lindblad master equation
@@ Acknowledgements.
$|\phi(t+\delta t)\rangle=e^{-iH\delta t}\mathcal{N}\left\{\begin{matrix}(1)\left\{\begin{matrix}\delta p_{1}:&L_{1}\\ 1-\delta p_{1}:&\exp(-\frac{1}{2}\gamma_{1}L^{\dagger}_{1}L_{1}\delta t)\end{matrix}\right.\\ \vdots\\ (\mu)\left\{\begin{matrix}\delta p_{\mu}:&L_{\mu}\\ 1-\delta p_{\mu}:&\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)\end{matrix}\right.\\ \vdots\\ (M)\left\{\begin{matrix}\delta p_{M}:&L_{M}\\ 1-\delta p_{M}:&\exp(-\frac{1}{2}\gamma_{M}L^{\dagger}_{\mu}L_{M}\delta t)\end{matrix}\right.\\ \end{matrix}\ \ \ \right\}|\phi(t)\rangle.$ (S1)
@@
$\frac{(1-\eta_{\mu})\delta p_{\mu}}{1-\eta_{\mu}\delta p_{\mu}}\approx(1-\eta_{\mu})\delta p_{\mu},$ (S2)
@@
$\frac{1-\delta p_{\mu}}{1-\eta_{\mu}\delta p_{\mu}}\approx 1-(1-\eta_{\mu})\delta p_{\mu}.$ (S3)
@@
$|\phi(t+\delta t)\rangle=e^{-iH\delta t}\mathcal{N}\left\{\begin{matrix}(1)\left\{\begin{matrix}(1-\eta_{1})\delta p_{1}:&L_{1}\\ 1-(1-\eta_{1})\delta p_{1}:&\exp(-\frac{1}{2}\gamma_{1}L^{\dagger}_{1}L_{1}\delta t)\end{matrix}\right.\\ \vdots\\ (\mu)\left\{\begin{matrix}(1-\eta_{\mu})\delta p_{\mu}:&L_{\mu}\\ 1-(1-\eta_{\mu})\delta p_{\mu}:&\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)\end{matrix}\right.\\ \vdots\\ (M)\left\{\begin{matrix}(1-\eta_{M})\delta p_{M}:&L_{M}\\ 1-(1-\eta_{M})\delta p_{\mu}:&\exp(-\frac{1}{2}\gamma_{M}L^{\dagger}_{\mu}L_{M}\delta t)\end{matrix}\right.\\ \end{matrix}\ \ \ \right\}|\phi(t)\rangle.$ (S4)
@@
Making the first-order approximation of $\delta t$, the maximum probability state $|\phi_{eff}\rangle$ and its probability $P_{Max}$ in Eq. (S4) are
$P_{Max}=\prod_{\mu}[1-(1-\eta_{\mu})\delta p_{\mu}]\approx 1-\sum_{\mu}(1-\eta_{\mu})\delta p_{\mu},$ (S5)
$\begin{split}|\phi_{eff}\rangle&=e^{-iH\delta t}\mathcal{N}\left(\prod_{\mu}\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)\right)|\phi(t)\rangle\\ &\approx\mathcal{N}\exp(-iH_{eff}\delta t)|\phi(t)\rangle=\frac{1}{\sqrt{1-\sum_{\mu}\delta p_{\mu}}}\exp(-iH_{eff}\delta t)|\phi(t)\rangle,\\ \end{split}$ (S6)
@@
$P_{\mu}=(1-\eta_{\mu})\delta p_{\mu}\prod_{j\neq\mu}[1-(1-\eta_{j})\delta p_{j}]\approx(1-\eta_{\mu})\delta p_{\mu},$ (S7)
$\begin{split}|\phi_{\mu}\rangle&=e^{-iH\delta t}\mathcal{N}\left(\prod_{j\neq\mu}\exp(-\frac{1}{2}\gamma_{j}L^{\dagger}_{j}L_{j}\delta t)\right)L_{\mu}|\phi(t)\rangle\\ &\approx\mathcal{N}\exp(-iH\delta t)L_{\mu}|\phi(t)\rangle=\frac{1}{\sqrt{<L_{\mu}^{{\dagger}}L_{\mu}>}}\exp(-iH\delta t)L_{\mu}|\phi(t)\rangle.\\ \end{split}$ (S8)
@@
$\begin{split}\frac{d}{dt}\rho&=\lim_{\delta t\to 0}\frac{1}{\delta t}(|\phi(t+\delta t)\rangle\langle\phi(t+\delta t)|-|\phi(t)\rangle\langle\phi(t)|)\\ &=\lim_{\delta t\to 0}\frac{1}{\delta t}\left(P_{Max}|\phi_{eff}\rangle\langle\phi_{eff}|+\left(\sum_{\mu}P_{\mu}|\phi_{\mu}\rangle\langle\phi_{\mu}|\right)-|\phi(t)\rangle\langle\phi(t)|\right)\\ &=\lim_{\delta t\to 0}\frac{1}{\delta t}\left(\frac{1-\sum_{\mu}(1-\eta_{\mu})\delta p_{\mu}}{1-\sum_{\mu}\delta p_{\mu}}e^{-iH_{eff}\delta t}\rho e^{iH^{\dagger}_{eff}\delta t}+\sum_{\mu}\left(\frac{(1-\eta_{\mu})\delta p_{\mu}}{<L_{\mu}^{{\dagger}}L_{\mu}>}e^{-iH\delta t}L_{\mu}\rho L^{\dagger}_{\mu}e^{iH\delta t}\right)-\rho\right)\\ &=\lim_{\delta t\to 0}\frac{1}{\delta t}\left((1+\sum_{\mu}\eta_{\mu}\delta p_{\mu})[\rho-i\delta t(H_{eff}\rho-\rho H^{\dagger}_{eff})]+\sum_{\mu}\left((1-\eta_{\mu})\gamma_{\mu}\delta tL_{\mu}\rho L^{\dagger}_{\mu}\right)-\rho\right)\\ &=-i[H,\rho]+\sum_{\mu}\gamma_{\mu}\left(-\frac{1}{2}\{L_{\mu}^{{\dagger}}L_{\mu},\rho\}+(1-\eta_{\mu})L_{\mu}\rho L_{\mu}^{{\dagger}}+\eta_{\mu}<L_{\mu}^{{\dagger}}L_{\mu}>\rho\right).\\ \end{split}$ (S9)
S2. Vectorization treatment of the two-level atom.

### v3

@@ Abstract
I Introduction
The Lindblad master equation (LME) is one of the most successful dynamics equations in the quantum open system Lindblad1976 ; Gorini1976 ; Gardiner1985 ; Daniel2020 . Its success is not only attributed to the effective description of the quantum system weakly coupling to the environment but also stems from its concise mathematical structure with good properties such as trace-preserving, trace-preserving, and complete positivity. From the perspective of quantum trajectory theory Dalibard1992 ; Dum1992 ; Gardiner1992 ; Klaus1993 ; Plenio1998 ; Daley2014 , the LME can be understood as an average of many quantum trajectories consisting of the evolution of non-Hermitian Hamiltonian (ENHH) and random quantum jumps. Due to the constraint by LME, ENHH and the quantum jumps are required to reach a balance in the average sense of all trajectories to keep some good properties like trace-preserving.
@@ I Introduction
II Nonlinear Lindblad master equation
Considering a system with Hamiltonian $H$ coupling to the environment by $M$ jump operators $L_{\mu}$ with strength $\gamma_{\mu}$, its density matrix $\rho$ is governed by LME:
@@ II Nonlinear Lindblad master equation
where $\mathcal{L}$ is the Liouvillian superoperator and Planck constant $\hbar$ is set to unity throughout this paper. We monitor every jump by a $\eta_{\mu}$-efficiency detector, $\eta_{\mu}\in[0,1]$, as shown in Fig. 1(a). The core concern of the present work is how $\rho$ evolves after discarding the data along with detectors clicking. When $\eta_{\mu}=0$, the detector does not work, meaning all quantum trajectories are retained, and the dynamics is fully described by the LME. Conversely, for an ideal detector ($\eta_{\mu}=1$), every quantum jump is perfectly monitored, and all trajectories containing jumps are discarded, resulting in postselection of purely ENHH. For a finite efficiency detector, the quantum trajectories will be probabilistically discarded, so the postselected process is the Lindbladian dynamics with loss of part of quantum jumps.
@@
we get the NLME (see Appendix  A for details):
@@
The simplest model for Eq. (4) is a two-level atom with monitoring its spontaneous emission, as shown in Fig. 1(b), where $J$ is the coupling strength and $\gamma$ is the spontaneous emission rate of the excited state $|e\rangle$. The detector of emission photons has a finite efficiency $\eta$. This setup is also discussed in Ref. Lewalle2023 . In our theoretical study, the observed data of atomic population is discarded when emission photons are detected simultaneously. With the remaining data, the evolution of atomic system is rebuilt by
@@
A vectorization treatment of Eq. (9) is given in Appendix  B. We further compare the results from this method with results from two kinds of quantum trajectory methods in Appendix  D to confirm the validity of NLME.
III Properties and classification
The NLME given by Eq. (4) provides a convenient method to study the postselection dynamics. Before going to the study of concrete systems, we discuss some specific features of the NLME. Firstly, we show some properties of NLME.
@@ III Properties and classification
IV An example in the trivial class
Considering a homogeneous fermion chain in periodic boundary condition under $\eta$-efficiency continuous measurements of particle number in every site. The postselection dynamics is described by
@@ IV An example in the trivial class
V Postselected skin effect
Now we study a model in the nontrivial class, which exhibits skin effect induced by postselection. Considering a $L$-site free fermion chain under open boundary condition monitored by $L+1$ detectors with unified efficiency $\eta$ as shown in Fig. 2(a), the corresponding NLME is given by
@@ V Postselected skin effect
where $H=J\sum_{l=1}^{L-1}a_{l}^{\dagger}a_{l+1}+h.c.$, $\gamma_{l}=\gamma$, $L_{l}=d_{l}^{\dagger}d_{l}$ with $d_{l}=(a_{l}+ia_{l+1})/\sqrt{2}$ for $l=1\sim L-1$, $\gamma_{0}=\gamma_{L}=\gamma/2$, $L_{0}=\hat{n}_{1}$ and $L_{L}=\hat{n}_{L}$. For simplicity, we set $\gamma\leq J=1$. We note that the jump operator $L_{l}$ originates from the Ref. Wang2022 , where a biased steady state is induced through the combination between $L_{l}$ and some feedback unitary operators. In this work, we demonstrate that a biased steady state can also be achieved by replacing the feedback unitary operators with postselection.
Figure 2: Postselected skin effect. (a) Illustration of the model with postselected skin effect. (b) Steady-state distributions of the half-filled chain with 50 (red star), 100 (blue circle), and 150 (black triangle) sites. The green line is the fitting function Eq. (22) with $\beta=0.04$. (c) Evolution of half-chain entanglement entropy defined by Eq. (25) in the half-filled chain with 50 (red line), 100 (blue line), and 150 (black line) sites. In (b) and (c), $\gamma=0.4$, $\eta=0.6$ and the initial state is $|1010\cdots\rangle$. The results are an average of $60$ quantum trajectories by the Gaussian state method and the Monte Carlo wave-function method (see Appendix C) with time step $\delta t=0.005$.
@@
VI Zeno effect versus skin effect
The NLME of postselected skin effect formally is same as the Hatano-Nelson model in open system. The effective Hamiltonian $H_{eff}(\eta\gamma)$ in Eq. (23a) leads the non-Hermitian skin effect Yao2018 ; Kunst2018 ; CHLee2019 , while the environmental effect $\mathcal{L}_{M}$ in Eq. (23b) can be understood as continuous measurements, which induces the quantum Zeno effect Misra1997 . Considering that both of Zeno effect and skin effect can influence the entangled entropy growth, the TAEE $\overline{S}_{L}(x_{C},x_{C}+\Delta)$ will be a good indicator to reflect the two effects’ competition and cooperation. The TAEE is defined as
@@ VI Zeno effect versus skin effect
VII Distribution and scaling of entangled entropy
The postselected skin effect profoundly affects the distribution and scaling behaviors of TAEE by the tanh-fitting steady state, where the particle number only changes rapidly in the middle of the chain. Near the ends of a long chain, the local site is nearly fully occupied ($n_{x}\approx 1$) or near vacuum ($n_{x}\approx 0$), which contributes little to entangled entropy. This makes the entropy distribution $\overline{S}_{L}(x_{C},x_{C}+\Delta)$ exhibit a two-segment structure—a nearly algebraic growth with a small $\Delta$ followed by saturation for large $\Delta$, as shown in Fig. 4(a).
Figure 4: Distribution and scaling of steady-state entangled entropy. (a) Steady-state entanglement entropy $\overline{S}_{L}(x_{C},x_{C}+\Delta)$ as a function of the bipartition size $\Delta$ in a half-filled $150$-site chain with $\eta=0.6$ and $\gamma$ from 00 to $1$ in the log-log plot. (b) Half-chain entanglement entropy $\overline{S}_{L}(x_{C},x_{C}+L/2)$ scaling with the half-filled chain length $L$ in the log-log plot. The entropy has a two-segment scaling behavior: rapid growth in small $L$ and saturation in large $L$. With decreasing $\gamma$, the rapid growth segment becomes longer. Especially in $\gamma=0$, the growth segment increases to the whole chain, which implies a volume law of entropy in the non-dissipative free systems.
The steady state’s scaling behavior of the half-chain TAEE $\overline{S}_{L}(x_{C},x_{C}+L/2)$ also presents a two-segment structure. When increasing $L$, the population distribution of the whole chain becomes a larger part of the tanh function Eq. (22), as shown in Fig. 4(b), and meanwhile the end-site particle number $n_{L}$ decreases rapidly from $0.5$ and then approaches slowly to 00. This leads to $\overline{S}_{L}(x_{C},x_{C}+L/2)$ having a nearly algebraic growth in small $L$ and saturation in large $L$, as shown in Fig. 4(b).
VIII Summary and discussion
In summary, we proposed a NLME to describe the normalized postselection dynamics of Lindblad master equations with loss of quantum jumps. It continuously interpolates between LME and ENHH by reducing quantum jumps. With the advanced analytical ability of NLME, we answer the question of how it influences the dynamics of LME by losing quantum jumps. We classify the influences into two classes. In trivial class, discarding jumps is completely equivalent to weakening the coupling strength of the environment. While in the nontrivial class, it makes a difference by breaking the balance of LME between the effective non-Hermitain Hamiltonian and quantum jumps. We show a new phenomenon called the postselected skin effect, where the steady state exhibits a special biased particle profile by discarding a few quantum jumps. Moreover, we show the coexistence and competition between Zeno effect and non-Hermitian skin effect in the postselected skin effect by capturing the characteristics of the TAEE.
@@ VIII Summary and discussion
Acknowledgements
We thank Y.-P. Wang for helpful discussions on the Gaussian-state method. This work is supported by National Key Research and Development Program of China (Grant No. 2023YFA1406704 and 2021YFA1402104), the NSFC under Grants No. 12174436 and No. T2121001, and the Strategic Priority Research Program of Chinese Academy of Sciences under Grant No. XDB33000000.
Appendix A The derivation of the nonlinear Lindblad master equation
@@
$\begin{split}|\phi(t+\delta t)\rangle=&e^{-iH\delta t}\\ &\mathcal{N}\left\{\begin{matrix}(1)\left\{\begin{matrix}\delta p_{1}:&L_{1}\\ 1-\delta p_{1}:&\exp(-\frac{1}{2}\gamma_{1}L^{\dagger}_{1}L_{1}\delta t)\end{matrix}\right.\\ \vdots\\ (\mu)\left\{\begin{matrix}\delta p_{\mu}:&L_{\mu}\\ 1-\delta p_{\mu}:&\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)\end{matrix}\right.\\ \vdots\\ (M)\left\{\begin{matrix}\delta p_{M}:&L_{M}\\ 1-\delta p_{M}:&\exp(-\frac{1}{2}\gamma_{M}L^{\dagger}_{\mu}L_{M}\delta t)\end{matrix}\right.\\ \end{matrix}\ \ \ \right\}|\phi(t)\rangle.\end{split}$ (27)
@@
$\frac{(1-\eta_{\mu})\delta p_{\mu}}{1-\eta_{\mu}\delta p_{\mu}}\approx(1-\eta_{\mu})\delta p_{\mu},$ (28)
@@
$\frac{1-\delta p_{\mu}}{1-\eta_{\mu}\delta p_{\mu}}\approx 1-(1-\eta_{\mu})\delta p_{\mu}.$ (29)
@@
$\begin{split}&|\phi(t+\delta t)\rangle=e^{-iH\delta t}\\ &\mathcal{N}\left\{\begin{matrix}(1)\left\{\begin{matrix}(1-\eta_{1})\delta p_{1}:&L_{1}\\ 1-(1-\eta_{1})\delta p_{1}:&\exp(-\frac{1}{2}\gamma_{1}L^{\dagger}_{1}L_{1}\delta t)\end{matrix}\right.\\ \vdots\\ (\mu)\left\{\begin{matrix}(1-\eta_{\mu})\delta p_{\mu}:&L_{\mu}\\ 1-(1-\eta_{\mu})\delta p_{\mu}:&\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)\end{matrix}\right.\\ \vdots\\ (M)\left\{\begin{matrix}(1-\eta_{M})\delta p_{M}:&L_{M}\\ 1-(1-\eta_{M})\delta p_{M}:&\exp(-\frac{1}{2}\gamma_{M}L^{\dagger}_{\mu}L_{M}\delta t)\end{matrix}\right.\\ \end{matrix}\ \ \ \right\}|\phi(t)\rangle.\end{split}$ (30)
@@
Making the first-order approximation of $\delta t$, the maximum probability state $|\phi_{eff}\rangle$ and its probability $P_{Max}$ in Eq. (30) are
$P_{Max}=\prod_{\mu}[1-(1-\eta_{\mu})\delta p_{\mu}]\approx 1-\sum_{\mu}(1-\eta_{\mu})\delta p_{\mu},$ (31)
$\begin{split}|\phi_{eff}\rangle&=e^{-iH\delta t}\mathcal{N}\left(\prod_{\mu}\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)\right)|\phi(t)\rangle\\ &\approx\mathcal{N}\exp(-iH_{eff}\delta t)|\phi(t)\rangle\\ &=\frac{1}{\sqrt{1-\sum_{\mu}\delta p_{\mu}}}\exp(-iH_{eff}\delta t)|\phi(t)\rangle,\\ \end{split}$ (32)
@@
$P_{\mu}=(1-\eta_{\mu})\delta p_{\mu}\prod_{j\neq\mu}[1-(1-\eta_{j})\delta p_{j}]\approx(1-\eta_{\mu})\delta p_{\mu},$ (33)
$\begin{split}|\phi_{\mu}\rangle&=e^{-iH\delta t}\mathcal{N}\left(\prod_{j\neq\mu}\exp(-\frac{1}{2}\gamma_{j}L^{\dagger}_{j}L_{j}\delta t)\right)L_{\mu}|\phi(t)\rangle\\ &\approx\mathcal{N}\exp(-iH\delta t)L_{\mu}|\phi(t)\rangle\\ &=\frac{1}{\sqrt{<L_{\mu}^{{\dagger}}L_{\mu}>}}\exp(-iH\delta t)L_{\mu}|\phi(t)\rangle.\\ \end{split}$ (34)
@@
$\begin{split}\frac{d}{dt}\rho&=\lim_{\delta t\to 0}\frac{1}{\delta t}(|\phi(t+\delta t)\rangle\langle\phi(t+\delta t)|-|\phi(t)\rangle\langle\phi(t)|)\\ &=\lim_{\delta t\to 0}\frac{1}{\delta t}\left(P_{Max}|\phi_{eff}\rangle\langle\phi_{eff}|+\left(\sum_{\mu}P_{\mu}|\phi_{\mu}\rangle\langle\phi_{\mu}|\right)-|\phi(t)\rangle\langle\phi(t)|\right)\\ &=\lim_{\delta t\to 0}\frac{1}{\delta t}\left(\frac{1-\sum_{\mu}(1-\eta_{\mu})\delta p_{\mu}}{1-\sum_{\mu}\delta p_{\mu}}e^{-iH_{eff}\delta t}\rho e^{iH^{\dagger}_{eff}\delta t}+\sum_{\mu}\left(\frac{(1-\eta_{\mu})\delta p_{\mu}}{<L_{\mu}^{{\dagger}}L_{\mu}>}e^{-iH\delta t}L_{\mu}\rho L^{\dagger}_{\mu}e^{iH\delta t}\right)-\rho\right)\\ &=\lim_{\delta t\to 0}\frac{1}{\delta t}\left((1+\sum_{\mu}\eta_{\mu}\delta p_{\mu})[\rho-i\delta t(H_{eff}\rho-\rho H^{\dagger}_{eff})]+\sum_{\mu}\left((1-\eta_{\mu})\gamma_{\mu}\delta tL_{\mu}\rho L^{\dagger}_{\mu}\right)-\rho\right)\\ &=-i[H,\rho]+\sum_{\mu}\gamma_{\mu}\left(-\frac{1}{2}\{L_{\mu}^{{\dagger}}L_{\mu},\rho\}+(1-\eta_{\mu})L_{\mu}\rho L_{\mu}^{{\dagger}}+\eta_{\mu}<L_{\mu}^{{\dagger}}L_{\mu}>\rho\right).\\ \end{split}$ (35)
Appendix B Vectorization treatment of the two-level atom